### v1

@@ 3.3.1 Overview
The concept of environmental sustainability in AI systems emphasizes a balance between the environment, society, and economy [14], ensuring that AI innovations contribute positively to each area without causing adverse effects on the environment. Researchers recognize the significant carbon footprint and energy demands of extensive IoT networks and cloud/edge communications, advocating for an architectural and deployment approach that minimizes environmental impact [16]. The environmental impact of training a single extensive machine learning model, like Meena, can be equated to driving an average passenger vehicle for 242,231 miles [152]. However, this represents merely one facet of the broader picture. The environmental cost not only exists in the AI model training process but also in the AI system life cycle. As depicted in Figure 4, we explore potential environmental costs at each stage of the AI development pipeline. After that, we summarize the strategies applied to optimize the cost at each stage. Finally, we analyze the consequences of constructing an environmental sustainability AI system, focusing on its implications for safety, transparency, non-discrimination, and traceability.
@@
-Data Acquisition and Preparation: In this phase, energy consumption is predominantly driven by two processes: (1) the loading and validation split of large datasets, and (2) data preprocessing and feature extraction. The former necessitates downloading and storing training data on the model training server, a process that, although essential, is relatively less energy-intensive. Bouza et al. [153] observed that loading a 6 GB Imagenet validation split accounted for merely 0.5% of the total energy consumption compared to model training. The latter process involves feature extraction and the weighting of individual features to assess the model’s training efficacy, requiring significant computational resources. This stage is marked by exploring a wide array of machine-learning concepts on a large scale. Wu et al. [16] reported that the energy expenditure for this process at Meta was half that of the model training phase.
@@
Green-Algorithm101010www.green-algorithms.org Energy: avg 12W/core if UNK; TDP. Usage Factor: 100% usage if UNK. Energy: TDP; nvidia-smi; or avg 200W/GPU. Usage Factor: 100% usage if UNK Energy: 0.3725W/GB of memory available. Usage Factor: all available/requested memory. Location: not restricted. Static data: Carbon Foot-print; Electricity Maps. Default: 475gCO2eq/kWh
Codecarbon [^11] https://github.com/responsibleproblemsolving/energy-usage Energy: RAPL files or Power Gadget; TDP. Usage Factor: avg value 50% Energy: pynvml library (NVIDIA GPUs only). Usage Factor: whole machine consumption Energy: 0.3725W/GB of memory available. Usage Factor: allocated memory by ’process’ US & Canada: regional emissions per unit of power consumed; Other Countries: avg energy mix from Global Petrol Prices. Default: 475gCO2eq/kWh
@@
CarbonTracker [^13] https://github.com/lfwa/carbontracker Energy: RAPL files. Usage Factor: whole machine RAPL value Energy: pynvml library (NVIDIA GPUs only). Usage Factor: whole machine consumption Energy: use RAPL files. Usage Factor: used by ’process’ & others Real-time data: Energi Data Service (Denmark); Carbon Intensity API (Great Britain). Static data: carbon-intensities.csv. Default: 475gCO2eq/kWh
@@
Cumulator [^16] https://github.com/EPFLiGHT/cumulator Energy: avg 250W/core if UNK; TDP. Usage Factor: None Energy: avg 250W/GPU if UNK; TDP. Usage Factor: max load None Static data: Electricity Maps. Default: 447gCO2eq/kWh
@@
-Model Training: Upon selecting an appropriate training solution, the AI model undergoes training using a more comprehensive set of production data, which is not only more current but also larger in volume and enriched with features. The carbon emissions associated with model training can be divided into two types: offline and online training emissions. Offline training leverages historical data for model training, whereas online training dynamically updates the model parameters using the latest data. Wu et al. noted that, at Meta, the carbon emissions from online training were significantly lower than those from offline training [16].
@@
Additionally, the selection of batch sizes and epoch configurations influences energy consumption. Selecting an optimal batch size represents a balance between energy consumption and runtime [153], with larger batch sizes offering faster runtime and reduced energy usage. Nonetheless, excessively large batch sizes can lead to inefficient energy consumption due to low usage rates. Therefore, the ideal batch size is the maximum size that ensures full utilization of all GPUs and CPUs. Similarly, epoch configuration compromises the model’s quality and runtime. The duration of epochs and their associated energy consumption (or carbon emissions) remain constant [154]. A larger number of epochs extends the runtime, while also enhancing the likelihood of achieving model convergence. The model structure configuration directly impacts the amount of the model’s parameters.
@@
Beyond the previously mentioned factors, checkpointing exerts a negligible effect, and the impact of the number of parameters on energy use and carbon emissions is uncertain. Bouza et al. [153] identified an inconsequential difference in energy consumption between experiments conducted with a single checkpoint versus ten checkpoints. Typically, a larger number of parameters suggests extensive computational efforts during training. Nevertheless, actual energy consumption varies based on the specific model setup and training infrastructure employed. For example, training a Switch Transformer model, which boasts 1.5 trillion parameters [155], results in considerably lower carbon emissions compared to the GPT-3 model, which contains 750 billion parameters [156].
@@
-AI System Manufacturing and Deployment: The manufacturing and deployment of AI systems significantly affect carbon emissions and energy consumption. The impacts related to AI system manufacturing and deployment include the model’s implementation program, the quantity and capabilities of processors running the program, the efficiency of data centers in power delivery and cooling, and the mix of energy sources (e.g., renewable, gas, coal) [152]. For instance, Meta’s Transformer-based Universal Language Model for text translation at the inference stage served a lot of production traffic resulting in large energy consumption [157]. Bouza et al. [153] also highlight the impact of infrastructure size on power consumption through comparative experiments conducted on different platforms. For instance, training the Denoiser model on the larger Gemini-1 infrastructure (Grid5000 server) was completed in 2 hours, whereas it took 3 hours and 16 minutes on the smaller Rosenblatt infrastructure (MAP5 server). Furthermore, the CPU usage factor was significantly lower on Grid5000 (16%) compared to MAP5 (39%), and a similar trend was observed for GPU usage, with 14.3% on Grid5000 and 54% on MAP5, indicating higher efficiency in larger infrastructures. The energy-intensive nature of producing advanced computing components, coupled with the electricity demands of data centers hosting AI applications, underscores the environmental implications of AI’s widespread adoption.
@@
A denial-of-service (DoS) attack at the AI system deployment stage can generate sponge examples, significantly increasing energy consumption and the model’s inference run time [158]. This attack exploits the energy-latency gap vulnerability in AI systems, where different inputs of the same size can cause a DNN to consume varying amounts of time and energy. This disparity is due to optimization strategies in hardware and AI algorithms. For instance, custom and semi-custom hardware typically leverage data sparsity and low-precision computations for DNN inference, reducing both arithmetic complexity and DRAM traffic to achieve better power efficiency [159, 160, 161]. However, attackers can craft sponge examples by manipulating the token size of an NLP model’s input and output sentences and the size of embedding spaces, leading to a substantial increase in algorithmic complexity and energy consumption. A proposed defense is to set a worst-case performance bound for some models and monitor their usage [158].
@@
Table 5 summarized a list of tools [153] used to measure CPU, GPU, and memory consumption and the emission efficiency related to the AI system’s manufacturing and deployment. Users can monitor their AI system’s environmental cost in its development and deployment to determine which stage needs to be improved to achieve environmental sustainability.
@@
-Environmental cost optimizations: In the realm of AI, environmental cost optimization encompasses several pivotal areas, including algorithmic enhancements, processor improvements, data center efficiency, and the strategic mix of energy sources. A notable advancement in algorithm development is illustrated by the Evolved Transformer (Medium) model, which through neural architecture search, operates with significantly lower computational demands-1.6 times fewer FLOPS and reduced processing time by 1.1 to 1.3 times compared to its predecessor, the Transformer (Big), while simultaneously achieving a marginal increase in accuracy [152]. This progress underscores the potential for algorithmic and program improvements to contribute directly to environmental sustainability by reducing the energy footprint of AI operations.
@@
Further optimizations are evident in the development of specialized hardware for deep learning, aimed at enhancing the cost-performance ratio. The focus on Total Cost of Ownership (TCO), which includes operational costs like electricity consumption and the capital expenses associated with computing infrastructure, highlights the relationship between power usage and financial expenditure. Patterson et al. [152] shows that enhancements in performance per TCO also yield benefits in performance per watt, leading to financial savings and a reduction in carbon emissions. Moreover, cloud data centers exhibit significantly higher energy efficiency compared to traditional enterprise data centers, attributed to factors such as improved server utilization. This efficiency has mitigated the anticipated increase in energy consumption attributed to data center operations. Additionally, the strategic deployment of cloud computing facilities in locations with cleaner energy grids or where clean energy can be directly procured (e.g., Finland and Iowa) leverages the efficiency of transmitting information through optical fibers over long distances, further diminishing the environmental impact of AI systems. This holistic approach to environmental cost optimization in AI system manufacturing and deployment highlights the synergy between technological advancements and strategic operational decisions in reducing the carbon footprint and energy consumption of AI technologies.
@@ 3.3.2 Environment sustainability versus safety, transparency, non-discrimination, and traceability
[162] Model Fine-tuning A regression-focused profiling: find the various metrics’ trend with hardware deployment of NN Energy-efficient configuration for model development on the target device Safety (+), Non-discrimination (+)
@@
[163] A super-fine-tuning apply a dynamic learning rate method Fast learning with regular change learning rate end in low energy cost
[164], [165] Last-layer fine-tuning
@@
[166], [167], [168] Transfer Learning Adaptive or layer-wise fine-tuning: select layer to be fine-tuned Fewer layers fine-tuned reduce memory footprint & computational costs Non-discrimination (+/-), transparency (+)
[169], [170] Parameter-based learning share parameters between models
[171], [172] Relation-based learning establish effective model mapping
@@
[173], [174] Instance-based learning adapts data transferring knowledge Negative impact with training enlarged dataset
@@
[175] Phantom tree algorithm to measure the transfer cost Trade-off between memory and computational cost and model performance
@@
[100], [176] Machine Unlearning Exact unlearning with sub model/checkpoints retrained to entirely eliminate the data trace Lower computational cost by avoiding the entire model retraining Non-discrimination (+), transparency (+)
@@
[104], [103] Approximate unlearning with a gradient-based deletion approach to limit training data’s impact
@@
[177] Retrieval-Augmented Generation (RAG) Pre-trained language models with information retrieval methods to elevate text creation Lower computational cost with the elimination of training expenses Non-discrimination (+)
@@
[178], [179] Model compression Pruning: Reduce model parameters that contribute little in training Memory- & energy-efficiency in model deployment and/or development with reduced model size Non-discrimination (-), transparency (+/-), traceability (-)
@@
[180], [181] Quantization: Reduce the number of bits used to represent each weight
@@
[182], [183] Low-rank factorization: Identify model’s redundant parameters with matrix & tensor decomposition
@@
[184], [185], [186] Knowledge distillation: Transfer knowledge from a large to a smaller model & ensure the validity
@@
Model fine-tuning: Model fine-tuning stands as a cornerstone technique in machine learning, streamlining the enhancement of an existing model to better align with specific tasks or data sets. As a cornerstone of transfer learning, this method allows for the seamless transfer of knowledge from one domain to a related one, facilitating the swift customization of pre-trained models for new tasks with limited labeled data [166]. This approach not only significantly boosts the accuracy of models for targeted applications but also underscores the efficiency of model fine-tuning in elevating overall model performance.
@@
Model fine-tuning offers an eco-friendly alternative to the traditional, resource-intensive process of training a model from the ground up using extensive datasets [167, 164, 165]. By utilizing a smaller training dataset, fine-tuning significantly reduces computational and storage requirements during the Data Acquisition and Preparation phase. This reduction not only decreases the computational burden associated with loading, preprocessing, and extracting features from the data but also minimizes memory demands for data storage. In the Model Training phase, the smaller dataset and/or fewer layers to be trained facilitates quicker training times and manageable computational expenses. When applied within distributed or cloud-based systems, the minimized dataset size also leads to more manageable data communication costs, enhancing the overall efficiency and sustainability of the model development process.
@@
Additionally, this technique plays a pivotal role in creating environmentally sustainable AI by optimizing platform/application-level caching [16]. By leveraging pre-computed embeddings from existing models, fine-tuning improves power efficiency in developing new models for distinct tasks, thereby minimizing the environmental footprint of AI training processes. Mazumder et al. [162] suggested a profiling approach to identify the most efficient fine-tuning configuration in terms of energy consumption or latency, targeting the required accuracy. This method analyzes the energy/latency contours across various tinyML device platforms.
@@
Furthermore, model fine-tuning is instrumental in bolstering safety and fostering non-discrimination within AI development. Comprehensive fine-tuning can markedly diminish the threat of backdoor attacks, safeguarding the utility of AI systems while maintaining high safety standards [163, 187]. End-to-end fine-tuning effectively eradicates hidden vulnerabilities, showcasing its critical role in securing AI technologies. Moreover, adjusting the final layer of a model can enhance fairness, mitigating bias and promoting equality [188, 189, 190]. This harmonious integration of environmental sustainability, safety measures, and non-discriminatory practices through model fine-tuning exemplifies its comprehensive benefits in fostering the creation of ethical and responsible AI systems.
@@
Transfer learning: Transfer learning is advocated for scenarios where training data is either scarce or costly to acquire, enabling the use of different domains, tasks, and distributions for training and testing purposes [191, 192]. This approach, akin to model fine-tuning, promotes the development of eco-friendly AI by reducing energy consumption during the data acquisition, preparation, and model training stages. Unlike model fine-tuning, which focuses on preserving the pre-learned features, transfer learning seeks to adapt the pre-trained model’s weights, features, or architectural knowledge to a new task, requiring minimal further training [175]. In addition to benefits similar to those of fine-tuning, this discussion extends to examining the environmental advantages of transfer learning through its methods of transferring knowledge.
@@
To enhance model performance by facilitating the effective transfer of useful knowledge and preventing the transfer of detrimental knowledge during the Data Acquisition and Preparation phase. This effort is outlined in the work of Chen et al. [168] and further detailed by Niu et al. [174], who categorize the methodologies into four distinct types: instance-based, feature-based, parameter-based, and relation-based learning. Notably, parameter-based and relation-based methods are highlighted for their potential to support the development of AI systems in an environmentally sustainable manner.
@@
Instance-based and feature-based learning do not significantly contribute to the environmental sustainability of AI, while some specific methods may have negative impacts on it. Instance-based learning involves adapting data from the source domain for use in the target domain [173, 174], thereby enlarging the training set, which typically leads to training the model from scratch without any environmental benefits. Feature-based learning, while seeking domain-invariant features, can lead to a negative environmental impact. This type of learning often requires feature mapping techniques that transfer large amounts of data from the source to the target domain and involve adversarial learning and deep learning to extract high-level features. The latter methods can significantly amplify the computational burden during data acquisition and preparation. Wu et al. [175] introduced an innovative and cost-effective transfer learning framework, OPERA (Online Transfer using Phantom Tree for Real-Time Adaptation), to optimize the trade-off between accuracy improvement and computational expense. This framework employs the phantom tree algorithm, which assesses the complexity involved in constructing tree-based models, as a method to evaluate the costs associated with transfer learning.
@@
Conversely, learning about parameters and relations offers a more environmentally friendly approach. Parameter-based learning leverages shared parameters between models from the source and target domains, akin to model fine-tuning discussed above [169, 170]. Relation-based learning, through techniques like transferring a Markov Logic Network (MLN) or employing Relational Pathfinding (RPF), focuses on establishing effective mappings from the source model to the target domain [171, 172]. These strategies notably reduce the training time and data required to develop an accurate target domain model compared to building a model from scratch, as detailed by Mihalkova et al. [172]. This efficiency not only accelerates the learning process but also aligns with the principles of developing sustainable AI technologies by limiting the environmental footprint associated with extensive data processing and model training.

### v2

@@ 3.3.1 Overview
The concept of environmental sustainability in AI systems emphasizes a balance between the environment, society, and economy (?), ensuring that AI innovations contribute positively to each area without causing adverse effects on the environment. Researchers recognize the significant carbon footprint and energy demands of extensive IoT networks and cloud/edge communications, advocating for an architectural and deployment approach that minimizes environmental impact (?). The environmental impact of training a single extensive machine learning model, like Meena, can be equated to driving an average passenger vehicle for 242,231 miles (?). However, this represents merely one facet of the broader picture. The environmental cost not only exists in the AI model training process but also in the AI system life cycle. As depicted in Figure 4, we explore potential environmental costs at each stage of the AI development pipeline. After that, we summarize the strategies applied to optimize the cost at each stage. Finally, we analyze the consequences of constructing an environmental sustainability AI system, focusing on its implications for safety, transparency, non-discrimination, and traceability.
@@
-Data Acquisition and Preparation: In this phase, energy consumption is predominantly driven by two processes: (1) the loading and validation split of large datasets, and (2) data preprocessing and feature extraction. The former necessitates downloading and storing training data on the model training server, a process that, although essential, is relatively less energy-intensive. Bouza et al. (?) observed that loading a 6 GB Imagenet validation split accounted for merely 0.5% of the total energy consumption compared to model training. The latter process involves feature extraction and the weighting of individual features to assess the model’s training efficacy, requiring significant computational resources. This stage is marked by exploring a wide array of machine-learning concepts on a large scale. Wu et al. (?) reported that the energy expenditure for this process at Meta was half that of the model training phase.
@@
Green- Algori- thm101010www.green-algorithms.org Energy: avg 12W/core if UNK; TDP. Usage Factor: 100% usage if UNK. Energy: TDP; nvidia-smi; or avg 200W/GPU. Usage Factor: 100% usage if UNK Energy: 0.3725W/GB of memory available. Usage Factor: all available/requested memory. Location: not restricted. Static data: Carbon Foot-print; Electricity Maps. Default: 475gCO2eq/kWh
Code- carbon [^11] https://github.com/responsibleproblemsolving/energy-usage Energy: RAPL files or Power Gadget; TDP. Usage Factor: avg value 50% Energy: pynvml library (NVIDIA GPUs only). Usage Factor: whole machine consumption Energy: 0.3725W/GB of memory available. Usage Factor: allocated memory by ’process’ US & Canada: regional emissions per unit of power consumed; Other Countries: avg energy mix from Global Petrol Prices. Default: 475gCO2eq/kWh
@@
Carbon- Tracker [^13] https://github.com/lfwa/carbontracker Energy: RAPL files. Usage Factor: whole machine RAPL value Energy: pynvml library (NVIDIA GPUs only). Usage Factor: whole machine consumption Energy: use RAPL files. Usage Factor: used by ’process’ & others Real-time data: Energi Data Service (Denmark); Carbon Intensity API (Great Britain). Static data: carbon-intensities.csv. Default: 475gCO2eq/kWh
@@
Cumula- tor [^16] https://github.com/EPFLiGHT/cumulator Energy: avg 250W/core if UNK; TDP. Usage Factor: None Energy: avg 250W/GPU if UNK; TDP. Usage Factor: max load None Static data: Electricity Maps. Default: 447gCO2eq/kWh
@@
-Model Training: Upon selecting an appropriate training solution, the AI model undergoes training using a more comprehensive set of production data, which is not only more current but also larger in volume and enriched with features. The carbon emissions associated with model training can be divided into two types: offline and online training emissions. Offline training leverages historical data for model training, whereas online training dynamically updates the model parameters using the latest data. Wu et al. noted that, at Meta, the carbon emissions from online training were significantly lower than those from offline training (?).
@@
Additionally, the selection of batch sizes and epoch configurations influences energy consumption. Selecting an optimal batch size represents a balance between energy consumption and runtime (?), with larger batch sizes offering faster runtime and reduced energy usage. Nonetheless, excessively large batch sizes can lead to inefficient energy consumption due to low usage rates. Therefore, the ideal batch size is the maximum size that ensures full utilization of all GPUs and CPUs. Similarly, epoch configuration compromises the model’s quality and runtime. The duration of epochs and their associated energy consumption (or carbon emissions) remain constant (?). A larger number of epochs extends the runtime, while also enhancing the likelihood of achieving model convergence. The model structure configuration directly impacts the amount of the model’s parameters.
@@
Beyond the previously mentioned factors, checkpointing exerts a negligible effect, and the impact of the number of parameters on energy use and carbon emissions is uncertain. Bouza et al. (?) identified an inconsequential difference in energy consumption between experiments conducted with a single checkpoint versus ten checkpoints. Typically, a larger number of parameters suggests extensive computational efforts during training. Nevertheless, actual energy consumption varies based on the specific model setup and training infrastructure employed. For example, training a Switch Transformer model, which boasts 1.5 trillion parameters (?), results in considerably lower carbon emissions compared to the GPT-3 model, which contains 750 billion parameters (?).
@@
-AI System Manufacturing and Deployment: The manufacturing and deployment of AI systems significantly affect carbon emissions and energy consumption. The impacts related to AI system manufacturing and deployment include the model’s implementation program, the quantity and capabilities of processors running the program, the efficiency of data centers in power delivery and cooling, and the mix of energy sources (e.g., renewable, gas, coal) (?). For instance, Meta’s Transformer-based Universal Language Model for text translation at the inference stage served a lot of production traffic resulting in large energy consumption (?). Bouza et al. (?) also highlight the impact of infrastructure size on power consumption through comparative experiments conducted on different platforms. For instance, training the Denoiser model on the larger Gemini-1 infrastructure (Grid5000 server) was completed in 2 hours, whereas it took 3 hours and 16 minutes on the smaller Rosenblatt infrastructure (MAP5 server). Furthermore, the CPU usage factor was significantly lower on Grid5000 (16%) compared to MAP5 (39%), and a similar trend was observed for GPU usage, with 14.3% on Grid5000 and 54% on MAP5, indicating higher efficiency in larger infrastructures. The energy-intensive nature of producing advanced computing components, coupled with the electricity demands of data centers hosting AI applications, underscores the environmental implications of AI’s widespread adoption.
@@
A denial-of-service (DoS) attack at the AI system deployment stage can generate sponge examples, significantly increasing energy consumption and the model’s inference run time (?). This attack exploits the energy-latency gap vulnerability in AI systems, where different inputs of the same size can cause a DNN to consume varying amounts of time and energy. This disparity is due to optimization strategies in hardware and AI algorithms. For instance, custom and semi-custom hardware typically leverage data sparsity and low-precision computations for DNN inference, reducing both arithmetic complexity and DRAM traffic to achieve better power efficiency (?, ?, ?). However, attackers can craft sponge examples by manipulating the token size of an NLP model’s input and output sentences and the size of embedding spaces, leading to a substantial increase in algorithmic complexity and energy consumption. A proposed defense is to set a worst-case performance bound for some models and monitor their usage (?).
@@
Table 5 summarized a list of tools (?) used to measure CPU, GPU, and memory consumption and the emission efficiency related to the AI system’s manufacturing and deployment. Users can monitor their AI system’s environmental cost in its development and deployment to determine which stage needs to be improved to achieve environmental sustainability.
@@
-Environmental cost optimizations: In the realm of AI, environmental cost optimization encompasses several pivotal areas, including algorithmic enhancements, processor improvements, data center efficiency, and the strategic mix of energy sources. A notable advancement in algorithm development is illustrated by the Evolved Transformer (Medium) model, which through neural architecture search, operates with significantly lower computational demands-1.6 times fewer FLOPS and reduced processing time by 1.1 to 1.3 times compared to its predecessor, the Transformer (Big), while simultaneously achieving a marginal increase in accuracy (?). This progress underscores the potential for algorithmic and program improvements to contribute directly to environmental sustainability by reducing the energy footprint of AI operations.
@@
Further optimizations are evident in the development of specialized hardware for deep learning, aimed at enhancing the cost-performance ratio. The focus on Total Cost of Ownership (TCO), which includes operational costs like electricity consumption and the capital expenses associated with computing infrastructure, highlights the relationship between power usage and financial expenditure. Patterson et al. (?) shows that enhancements in performance per TCO also yield benefits in performance per watt, leading to financial savings and a reduction in carbon emissions. Moreover, cloud data centers exhibit significantly higher energy efficiency compared to traditional enterprise data centers, attributed to factors such as improved server utilization. This efficiency has mitigated the anticipated increase in energy consumption attributed to data center operations. Additionally, the strategic deployment of cloud computing facilities in locations with cleaner energy grids or where clean energy can be directly procured (e.g., Finland and Iowa) leverages the efficiency of transmitting information through optical fibers over long distances, further diminishing the environmental impact of AI systems. This holistic approach to environmental cost optimization in AI system manufacturing and deployment highlights the synergy between technological advancements and strategic operational decisions in reducing the carbon footprint and energy consumption of AI technologies.
@@ 3.3.2 Environment sustainability versus safety, transparency, non-discrimination, and traceability
(?) Model Fine-tuning A regression-focused profiling: find the various metrics’ trend with hardware deployment of NN Energy-efficient configuration for model development on the target device Safety (+), Non-discrimination (+)
@@
(?) A super-fine-tuning apply a dynamic learning rate method Fast learning with regular change learning rate end in low energy cost
(?, ?) Last-layer fine-tuning
@@
(?, ?, ?) Transfer Learning Adaptive or layer-wise fine-tuning: select layer to be fine-tuned Fewer layers fine-tuned reduce memory footprint & computational costs Non-discrimination (+/-), transparency (+)
(?, ?) Parameter-based learning share parameters between models
(?, ?) Relation-based learning establish effective model mapping
@@
(?, ?) Instance-based learning adapts data transferring knowledge Negative impact with training enlarged dataset
@@
(?) Phantom tree algorithm to measure the transfer cost Trade-off among memory, computational cost and model performance
@@
(?, ?) Machine Unlearning Exact unlearning with sub model/checkpoints retrained to entirely eliminate the data trace Lower computational cost by avoiding the entire model retraining Non-discrimination (+), transparency (+)
@@
(?, ?) Approximate unlearning with a gradient-based deletion approach to limit training data’s impact
@@
(?) RAG Pre-trained language models with information retrieval methods to elevate text creation Lower computational cost with the elimination of training expenses Non-discrimination (+)
@@
(?, ?) Model compression Pruning: Reduce model parameters that contribute little in training Memory- & energy-efficiency in model deployment and/or development with reduced model size Non-discrimination (-), transparency (+/-), traceability (-)
@@
(?, ?) Quantization: Reduce the number of bits used to represent each weight
@@
(?, ?) Low-rank factorization: Identify model’s redundant parameters with matrix & tensor decomposition
@@
(?, ?, ?) Knowledge distillation: Transfer knowledge from a large to a smaller model & ensure the validity
@@
Model fine-tuning: Model fine-tuning stands as a cornerstone technique in machine learning, streamlining the enhancement of an existing model to better align with specific tasks or data sets. As a cornerstone of transfer learning, this method allows for the seamless transfer of knowledge from one domain to a related one, facilitating the swift customization of pre-trained models for new tasks with limited labeled data (?). This approach not only significantly boosts the accuracy of models for targeted applications but also underscores the efficiency of model fine-tuning in elevating overall model performance.
@@
Model fine-tuning offers an eco-friendly alternative to the traditional, resource-intensive process of training a model from the ground up using extensive datasets (?, ?, ?). By utilizing a smaller training dataset, fine-tuning significantly reduces computational and storage requirements during the Data Acquisition and Preparation phase. This reduction not only decreases the computational burden associated with loading, preprocessing, and extracting features from the data but also minimizes memory demands for data storage. In the Model Training phase, the smaller dataset and/or fewer layers to be trained facilitates quicker training times and manageable computational expenses. When applied within distributed or cloud-based systems, the minimized dataset size also leads to more manageable data communication costs, enhancing the overall efficiency and sustainability of the model development process.
@@
Additionally, this technique plays a pivotal role in creating environmentally sustainable AI by optimizing platform/application-level caching (?). By leveraging pre-computed embeddings from existing models, fine-tuning improves power efficiency in developing new models for distinct tasks, thereby minimizing the environmental footprint of AI training processes. Mazumder et al. (?) suggested a profiling approach to identify the most efficient fine-tuning configuration in terms of energy consumption or latency, targeting the required accuracy. This method analyzes the energy/latency contours across various tinyML device platforms.
@@
Furthermore, model fine-tuning is instrumental in bolstering safety and fostering non-discrimination within AI development. Comprehensive fine-tuning can markedly diminish the threat of backdoor attacks, safeguarding the utility of AI systems while maintaining high safety standards . End-to-end fine-tuning effectively eradicates hidden vulnerabilities, showcasing its critical role in securing AI technologies. Moreover, adjusting the final layer of a model can enhance fairness, mitigating bias and promoting equality (?, ?, ?). This harmonious integration of environmental sustainability, safety measures, and non-discriminatory practices through model fine-tuning exemplifies its comprehensive benefits in fostering the creation of ethical and responsible AI systems.
@@
Transfer learning: Transfer learning is advocated for scenarios where training data is either scarce or costly to acquire, enabling the use of different domains, tasks, and distributions for training and testing purposes (?, ?). This approach, akin to model fine-tuning, promotes the development of eco-friendly AI by reducing energy consumption during the data acquisition, preparation, and model training stages. Unlike model fine-tuning, which focuses on preserving the pre-learned features, transfer learning seeks to adapt the pre-trained model’s weights, features, or architectural knowledge to a new task, requiring minimal further training (?). In addition to benefits similar to fine-tuning, this discussion examines the environmental advantages of transfer learning through its methods of transferring knowledge.
@@
To enhance model performance by facilitating the effective transfer of useful knowledge and preventing the transfer of detrimental knowledge during the Data Acquisition and Preparation phase. This effort is outlined in the work of Chen et al. (?) and further detailed by Niu et al. (?), who categorize the methodologies into four distinct types: instance-based, feature-based, parameter-based, and relation-based learning. Notably, parameter-based and relation-based methods are highlighted for their potential to support the development of AI systems in an environmentally sustainable manner.
@@
Instance-based and feature-based learning do not significantly contribute to the environmental sustainability of AI, while some specific methods may negatively impact it. Instance-based learning involves adapting data from the source domain for use in the target domain (?, ?), thereby enlarging the training set, which typically leads to training the model from scratch without any environmental benefits. Feature-based learning, while seeking domain-invariant features, can lead to a negative environmental impact. This type of learning often requires feature mapping techniques that transfer large amounts of data from the source to the target domain and involve adversarial learning and deep learning to extract high-level features. The latter methods can significantly amplify the computational burden during data acquisition and preparation. Wu et al. (?) introduced an innovative and cost-effective transfer learning framework, OPERA (Online Transfer using Phantom Tree for Real-Time Adaptation), to optimize the trade-off between accuracy improvement and computational expense. This framework employs the phantom tree algorithm, which assesses the complexity involved in constructing tree-based models, as a method to evaluate the costs associated with transfer learning.
@@
Conversely, learning about parameters and relations offers a more environmentally friendly approach. Parameter-based learning leverages shared parameters between models from the source and target domains, akin to model fine-tuning discussed above (?, ?). Relation-based learning, through techniques like transferring a Markov Logic Network (MLN) or employing Relational Pathfinding (RPF), focuses on establishing effective mappings from the source model to the target domain (?, ?). These strategies notably reduce the training time and data required to develop an accurate target domain model compared to building a model from scratch, as detailed by Mihalkova et al. (?). This efficiency not only accelerates the learning process but also aligns with the principles of developing sustainable AI technologies by limiting the environmental footprint associated with extensive data processing and model training.